OpenStudy (anonymous):

what's the inverse function of f(x)=x/(x-2)?

OpenStudy (holly00d1248):

Sorry dont know

OpenStudy (anonymous):

Let \(f(x) = y\)

OpenStudy (anonymous):

\[y = \frac{x}{x-2}\]

OpenStudy (anonymous):

Find \(x\) in terms of \(y\)..

OpenStudy (anonymous):

\[y(x-2) = x \\ xy - 2y = x \\ xy - x = 2y \\ x(y-1) = 2y \\ x = \frac{2y}{y-1}\]

OpenStudy (anonymous):

Now, just replace \(x\) with \(y\) and \(y\) with \(x\)..

OpenStudy (anonymous):

\[y = \frac{2x}{x-1}\]

OpenStudy (anonymous):

This is your inverse function..

OpenStudy (anonymous):

thank you so much!